An Everglades air boat travels 78 miles downstreat in the same amount of time it requires to move upstream a distance of 48 miles. the boats engines drive in still water at an average rate of 16 mph greater than the rate of the current. find the speed of the boat and the water current.
i tried
(16+c)t=48
(16+c)t=78
it didnt work. i cant figure this out please help!

Answer by stanbon(75887)

An Everglades air boat travels 78 miles downstreat in the same amount of time it requires to move upstream a distance of 48 miles. the boats engines drive in still water at an average rate of 16 mph greater than the rate of the current. find the speed of the boat and the water current.
i tried
(16+c)t=48
(16+c)t=78
it didnt work.
-------------------------------------------
Downstream DATA:
distance = 78 miles ; rate = (c+16+2 = (2c+16) mph ; time = 78/(2c+16) hrs.
----------------------
Upstream DATA:
distance = 48 miles ; rate = (16+c-c = 16 mph ; time = 48/16 = 3 hrs
-----------------------------------------------------------------
Equation:
time down = time up
78/(2c+16) = 3
78 = 6c + 48
6c = 30
c = 5 mph (speed of the current)
==================================
Cheers,
Stan H.